Circular No. 5950

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994D IN NGC 4526 A. Walker, Cerro Tololo Interamerican Observatory, reports the fol-

lowing preliminary data from CCD photometry obtained with the CTIO 0.9-m telescope on Mar. 10.2 UT: V = 13.66, B − V = −0.04, U − B = −0.4, V − I = +0.04.

Visual magnitude estimates: Mar. 9.932 UT, 13.3 (H. Dahle, Oslo, Norway); 10.885, 12.9 (O. Midtskogen, Tranby, Norway); 10.961, 12.5 (G. M. Hurst, Basingstoke, England); 11.548, 12.8 (Y. Sekino, Kanagawa, Japan); 12.012, 12.4 (Hurst).

GRB 940301 AND GRB 930704 J. M. Ryan, M. McConnell, R. M. Kippen, V. Schoenfelder, K. Ben-

nett, and W. Hermsen for COMPTEL; B. McNamara and T. E. Harrison, New Mexico State University; and H. Pedersen, Copenhagen University Observatory, write: “The COMPTEL team reports the potential repeat of a γ-ray burst that occurred on 1993 July 4.70075 UT. Based on available data, the COMPTEL position of GRB 930704 is α = 6^{h}43^{m}, δ = +62^{o}.5, with a 2σ error box covering 28 deg^{2}; the coordinates of GRB 940301 (IAUC 5943, 5944) are α = 6^{h}50^{m}, δ = +63^{o}.9, with a 2σ error box covering 3 deg^{2}. Without accounting for differential sky exposure, the probability of a 28deg^{2} overlap of burst positions is < 1 percent based on the low number of imaged COMPTEL bursts. Preliminary interplanetary annuli (provided by K. Hurley) intersect inside the 2σ error boxes of both bursts. If both bursts are assumed to come from the same source, the most likely position of the source, based solely on COMPTEL data, is approximately that of GRB 940301, by virtue of better statistics. For reference, at the center of the combined error boxes and annuli is the star SAOC 13977.”

PERIODIC COMET TEMPEL 1 (1993c) Total visual magnitude estimates (cf. IAUC 5940): Feb. 19.08 UT,

13.0 (G. Comello, Groningen, Netherlands, 0.18-m reflector); Mar. 3.30, 12.5 (R. Keen, Mt. Thorodin, CO, 0.32-m reflector); 5.26, 13.4 (G. W. Kronk, Troy, IL, 0.33-m reflector); 6.22, 12.8 (Kronk); 10.29, 12.6 (A. Hale, near Orogrande, New Mexico, 0.41-m reflector).

1994 March 12 Daniel W. E. Green